OpenStudy (anonymous):

Where is the center of the circle given by the equation (x + 4)2 + (y -3)2 = 4? Enter your answer as an ordered pair.

OpenStudy (cwrw238):

general form for a circle (x - a)^2 + (y - b)^2 = r^2 where (a,b) is the centre and r = radius

OpenStudy (anonymous):

ohh so my origin won;d be (4, 3)

OpenStudy (cwrw238):

now compare this with (x + 4)^2 + (y -3)^2 = 4

OpenStudy (cwrw238):

no, the 3 is correct but not the 4 - look at the sign in the first parenthesis

OpenStudy (anonymous):

ohh (-4. 3)

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

ok i feel dumb lol

OpenStudy (cwrw238):

(x + 4) = x- (-4))

OpenStudy (anonymous):

right :) thanks

OpenStudy (cwrw238):

yw